Calculate temperature in Fahrenheit if the temperature in celsius is -60C
F=(9/5)C+32.
I tried to solve the problem using .55 instead of 5/9 but my answer came out wrong.

Answer by nerdybill(7384)

Calculate temperature in Fahrenheit if the temperature in celsius is -60C
F=(9/5)C+32.
.
Simply plug and solve:
F=(9/5)C+32
F=(9/5)(-60)+32
F=(9)(-12)+32
F=-108+32
F=-76 (degrees Fahrenheit)